### v1

# Correlation between the charge radii difference in mirror partner nuclei and the symmetry energy slope

Abstract
Correlation between the charge radii difference of mirror partner nuclei $\Delta{R_{\mathrm{ch}}}$ and the slope parameter $L$ of symmetry energy has been built to ascertain the equation of state of isospin asymmetric nuclear matter. In this work, the influence of pairing correlations and isoscalar compression modulus on the $\Delta{R_{\mathrm{ch}}}$ are systematically investigated by Skyrme density functionals theory. The calculated results suggest that the linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$ is decreased by the surface pairing correlations. The slope parameter deduced from the difference of charge radii of mirror-pair nuclei 32Ar-32Si, 36Ca-36S, 38Ca-38Ar, and 54Ni-54Fe falls into the range of $L=42.57$-$50.64$ MeV, namely the rather soft equation of state of asymmetric nuclear matter. Besides, the range of slope parameter can also be influenced by the effective forces classified by various isoscalar incompressibility coefficients.
I INTRODUCTION
Nuclear symmetry energy can be generally employed to comprehend the underlying physical mechanism from the terrestrial nuclei to dense astrophysical events Steiner et al. (2005); Li et al. (2008); Roca-Maza and Paar (2018); Ji et al. (2019); Zhang et al. (2020); Xu et al. (2022a). So far, plenty of methods are undertaken to constrain the nuclear symmetry energy, such as the extracted neutron skin thickness (NST) Oyamatsu et al. (1998); Alex Brown (2000); Furnstahl (2002); Liu et al. (2011); Zhang and Chen (2013); Xu et al. (2020), properties of giant resonances Cao and Ma (2008); Cao et al. (2015); Roca-Maza et al. (2015, 2016); Cheng et al. (2023), quantities of the heavy-ion collision Lynch et al. (2009); Toro et al. (2010); Kumar et al. (2011); Xie et al. (2013); Feng and Jin (2010); Yu et al. (2020), etc. However, a unified density-dependence of symmetry energy is hardly determined due to the uncertainties suffering from different theoretical models. Thus more alternative observables should be urgently required in constraining the equation of state (EoS) of asymmetric nuclear matter. The difference of charge radii of mirror-pair nuclei has been proposed as a tentative probe to evaluate the slope parameter of symmetry energy $L$ Wang and Li (2013); Brown (2017). And then the precisely measured charge radii differences in mirror partner nuclei 32Ar-32Si, 36Ca-36S, 38Ca-38Ar, and 54Ni-54Fe are thereby used to validate the range of slope parameter $L$ Brown et al. (2020); Pineda et al. (2021); König et al. (2023).
Charge radii differences in mirror pair nuclei are strongly associated with the neutron-skin thickness of neutron-rich nuclei and with the slope parameter $L$ Wang and Li (2013); Brown (2017). As shown in Ref. Yang and Piekarewicz (2018), the calculations performed by relativistic energy density functionals (EDFs) further inspect the validity of this assumption. It points out that the difference in proton radii of mirror nuclei can be indeed employed to constrain the EoS of isospin asymmetry nuclear matter. This may provide an alternative approach to extract the content information about the neutron skin thickness as well Ding et al. (2023). Recent study suggests that mirror-difference in nuclear charge radii is proportional to the isospin asymmetry $I=(N-Z)/A$, where $N$ and $Z$ represent the neutron and proton number of a nucleus, respectively, and $A=N+Z$ is the corresponding mass number Novario et al. (2023). Here the same scenario can also be observed between the difference of binding energy per nucleon of mirror nuclei and the Coulomb asymmetry. Furthermore, combing the accurately detected charge radii differences in mirror partner nuclei, the upper limit range of symmetry energy slope has been determined at $L\approx 100$ MeV Bano et al. (2023).
Method mentioned above requires that more accurate charge radii data should be obtained through different nuclear models. With the improvement of experimental technique, high-precision charge radii of exotic nuclei can be possibly detected Yang et al. (2023a). Thus more charge radii data of nuclei far away from the $\beta$-stability line are compiled Angeli and Marinova (2013); Li et al. (2021). This provides a fundamental guideline for theoretical study to access the charge radii of mirror partner nuclei. A reliable description of nuclear charge radii is influenced by various mechanisms, such as shape deformation Lalazissis et al. (1996); Togashi et al. (2016); An et al. (2022a, 2023a), pairing correlations Reinhard and Nazarewicz (2017); An et al. (2020, 2022b), shell closure effects Nakada (2019); Perera et al. (2021), etc. As is well known, pairing correlations play an important role in describing the ground state properties of finite nuclei Meng (1998); Geng et al. (2003); An et al. (2018) or the quasiparticle resonant states Zhang and Qu (2020). Especially, it is mentioned that the radius of neutron-rich nuclei can also be influenced evidently by the pairing correlations Zhang et al. (2017).
As demonstrated in Ref. Reinhard and Nazarewicz (2022), pairing correlations should be taken into account in calculating the nuclear charge radii. Shown in Ref. Huang et al. (2023), the density energy functionals consisting of Skyrme and covariant models are undertaken to systematically investigate the correlations between the charge radii differences in mirror partner nuclei and slope parameter $L$. The calculated results suggest the linear correlation between the mirror difference in charge radii and slope parameter $L$ is further decreased by considering the pairing effects. In contrast, the correlation between the neutron-skin thickness of a neutron-rich nucleus and the proton radii difference of the corresponding mirror nuclei is mostly enhanced by pairing effects Ding et al. (2023). Meanwhile, it is also mentioned that the compression modulus of symmetric nuclear matter has an influence on determining the slope parameter $L$ An et al. (2023b).
The density-dependence of nuclear symmetry energy, which is mostly determined by the slope parameter $L$, is associated with the isovector-sensitive indicators in the EoS of isospin asymmetry systems. In Ref. Yoshida and Sagawa (2004), it suggests that the correlation between the isoscalar incompressibility $K$ and the NST is not clear. Further study demonstrates that the correlations between $K$ and the isovector parameters are generally weaker in comparison with the strong correlation between NST and symmetry energy coefficients Yoshida and Sagawa (2006); Xu et al. (2021). Actually, the nuclear matter properties recognizing the isoscalar and isovector counterparts are mutually correlated with the effective model parameters. This means the influence coming from the isoscalar sector is non-negligible in discussing the isovector properties Chatterjee et al. (2017); Mondal and Gulminelli (2023). As suggested in Refs. Piekarewicz (2004); Chen and Gu (2012); Kumar et al. (2021), the compression modulus of symmetric nuclear matter is sensitive to the density dependence of symmetry energy. Therefore, it is also essential to investigate the correlation between the difference of charge radii of mirror-partner nuclei and the slope parameter $L$ under various incompressibility coefficients $K$. In this work, the influence of pairing correlations and isoscalar compression modulus on the charge radii differences of mirror partner nuclei 32Ar-32Si, 36Ca-36S, 38Ca-38Ar, and 54Ni-54Fe are investigated by the spherical Skyrme EDFs. Furthermore, the range of slope parameter $L$ deriving from the charge radii differences of mirror-pair nuclei is discussed as well.
The structure of the paper is organized as the follows. Sec. II is devoted to presenting the theoretical framework briefly. In Sec. III, the numerical results and discussion are provided. Finally, a summary and outlook is given in Sec. IV.
II THEORETICAL FRAMEWORK
The sophisticated Skyrme-type energy density functional, which was expressed as an effective zero-range force between nucleons with density-dependent and momentum-dependent terms, has made remarkable success in describing various physical phenomena Bender et al. (2003); Pei et al. (2008, 2013); Wang et al. (2016); Zhang et al. (2017); Jiang and Wang (2018); Qu and Zhang (2019); Wu and Guo (2019); Jiang and Wang (2020a, b); Zhang and Qu (2020); Pei et al. (2014); Chai et al. (2020, 2020); Zhang et al. (2021); Chai et al. (2022); Wu et al. (2022); Huo et al. (2023). In general, the effective force sets have been determined by calibrating the properties of finite nuclei and symmetry nuclear matter at saturation density $\rho_{0}$($\approx 0.16~{}\mathrm{fm}^{-3}$). It is notably mentioned that the Skyrme EDF can give analytical expression of all variables characterizing infinite nuclear matter (see, e.g., Nakada (2003); Dutra et al. (2012) for details). In this work, the Skyrme-like effective interaction has been recalled as follows Köhler (1976); Chabanat et al. (1997, 1998),
$V(\mathbf{r}_{1},\mathbf{r}_{2})$ $=$ $t_{0}(1+x_{0}\mathbf{P}_{\sigma})\delta(\mathbf{r})$ (1)
$+\frac{1}{2}t_{1}(1+x_{1}\mathbf{P}_{\sigma})\left[\mathbf{P}^{\prime 2}\delta(\mathbf{r})+\delta(\mathbf{r})\mathbf{P}^{2}\right]$
$+t_{2}(1+x_{2}\mathbf{P}_{\sigma})\mathbf{P}^{\prime}\cdot\delta(\mathbf{r})\mathbf{P}$
$+\frac{1}{6}t_{3}(1+x_{3}\mathbf{P}_{\sigma})[\rho(\mathbf{R})]^{\alpha}\delta(\mathbf{r})$
$+\mathrm{i}W_{0}\mathbf{\sigma}\cdot\left[\mathbf{P}^{\prime}\times\delta(\mathbf{r})\mathbf{P}\right].$
Here, $\mathbf{r}=\mathbf{r}_{1}-\mathbf{r}_{2}$ and $\mathbf{R}=(\mathbf{r}_{1}+\mathbf{r}_{2})/2$ are related to the positions of two nucleons $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$, $\mathbf{P}=(\nabla_{1}-\nabla_{2})/2\mathrm{i}$ is the relative momentum operator and $\mathbf{P^{\prime}}$ is its complex conjugate acting on the left, and $\mathbf{P_{\sigma}}=(1+\vec{\sigma}_{1}\cdot\vec{\sigma}_{2})/2$ is the spin exchange operator that controls the relative strength of the $S=0$ and $S=1$ channels for a given term in the two-body interactions, with $\vec{\sigma}_{1(2)}$ being the Pauli matrices. The last term represents the spin-orbit force where $\sigma=\vec{\sigma}_{1}+\vec{\sigma}_{2}$. The quantities $\alpha$, $t_{i}$ and $x_{i}$ ($i=0$-3) represent the parameters of the Skyrme forces used in this work.
Solving the Schrödinger-like equations by the self-consistent iteration leads to the eigenenergies and wave functions of constituent nucleons. The pairing correlations can be generally treated either by the BCS method or by the Bogoliubov transformation Meng and Ring (1996); Meng (1998); Geng et al. (2003). In this work, spherical Skyrme Hartree-Fock-Bogoliubov (HFB) approach is utilized for all calculations Bennaceur and Dobaczewski (2005). The density-dependent zero-range pairing force is employed as follows Cao et al. (2012),
$V_{\mathrm{pair}}(\mathbf{r}_{1},\mathbf{r}_{2})=V_{0}\left[1-\eta\left(\frac{\rho(\mathbf{r})}{\rho_{0}}\right)\right]\delta(\mathbf{r}_{1}-\mathbf{r}_{2}),$ (2)
where $\rho(\mathbf{r})$ is the baryon density in coordinate space and $\rho_{0}=0.16~{}\mathrm{fm}^{-3}$ represents the nuclear saturation density. The value of $\eta$ is taken as $0.0$, $0.5$, or $1.0$ for volume-, mixed-, or surface-type pairing interactions, respectively. The quantity $V_{0}$ is adjusted by calibrating the empirical pairing gaps with three-point formula Cao et al. (2012); An et al. (2018), and the values of $V_{0}$ are 188.2 MeV fm3, 370.2 MeV fm3 and 509.6 MeV fm3 for the corresponding volume-, mixed-, and surface-type pairing interactions, respectively.
After achieving the convergence of the total binding energy, the root-mean square (rms) radii of the neutron and proton matter, and the neutron-skin thickness can be obtained naturally. The quantity of nuclear charge radius ($R_{\mathrm{ch}}$) can be calculated through (in units of fm2)
$R_{\mathrm{ch}}^{2}=\langle{r_{\mathrm{p}}^{2}}\rangle+0.64~{}\mathrm{fm^{2}}.$ (3)
The first term represents the charge density distributions of point-like protons and the second term is due to the finite size of protons Gambhir et al. (1990). For mirror pair nuclei, the difference in charge radii ($\Delta{R_{\mathrm{ch}}}$) can be obtained through the formula mentioned above.
Table 1: Parameters of the forces used in this work, where the parameter $\alpha=1/6$. Saturation properties of the Skyrme parametrization sets used in this work, such as the incompressibility coefficients $K$ (MeV), slope parameter $L$ (MeV) and symmetry energy $E_{\mathrm{sys}}$ (MeV) at saturation density $\rho_{0}$ (fm-3), are also shown.
   $K$    Sets     $t_{0}$      $t_{1}$     $t_{2}$     $t_{3}$     $x_{0}$     $x_{1}$     $x_{2}$     $x_{3}$     $W_{0}$     $L$     $E_{\mathrm{sys}}$
s3028 $-2461.903$ 472.082 $-530.239$ 13599.53 1.6830 $-0.3349$ $-1.0$ 2.6602 124.766 $-11.2262$ 28
s3030 $-2477.013$ 475.604 $-533.657$ 13709.09 1.1408 $-0.3365$ $-1.0$ 1.8302 117.482   22.8715 30
s3032 $-2491.849$ 486.614 $-579.661$ 13842.35 0.9962 $-0.2791$ $-1.0$ 1.5763 128.028   36.2246 32
$K\approx 230$ s3034 $-2503.455$ 489.112 $-591.648$ 13939.40 0.7308 $-0.2622$ $-1.0$ 1.1527 126.846   56.1442 34
s3036 $-2513.951$ 491.018 $-615.662$ 14045.31 0.5457 $-0.2210$ $-1.0$ 0.8437 128.528   71.5428 36
s3038 $-2524.594$ 494.060 $-635.510$ 14142.73 0.3550 $-0.1916$ $-1.0$ 0.5315 129.200   87.6155 38
s3040 $-2531.688$ 489.598 $-625.976$ 14203.48 0.1249 $-0.1940$ $-1.0$ 0.1662 122.045  106.0862 40
s4028 $-2296.534$ 515.586 $-336.464$ 11786.33 1.7714 $-0.8491$ $-1.0$ 3.1703 118.803    3.9774 28
s4030 $-2317.546$ 532.898 $-355.019$ 11907.07 1.2924 $-0.8449$ $-1.0$ 2.3824 118.263   34.0735 30
s4032 $-2321.087$ 520.235 $-491.504$ 12164.33 1.2626 $-0.5755$ $-1.0$ 2.1916 136.989   34.4283 32
$K\approx 240$ s4034 $-2329.794$ 510.932 $-423.708$ 12179.54 0.8546 $-0.6759$ $-1.0$ 1.5447 117.932   62.5884 34
s4036 $-2337.516$ 507.834 $-424.568$ 12253.22 0.7156 $-0.6664$ $-1.0$ 1.2882 118.414   75.6679 36
s4038 $-2355.765$ 523.051 $-465.145$ 12389.16 0.3790 $-0.6244$ $-1.0$ 0.7299 118.238   98.6522 38
s4040 $-2359.001$ 513.325 $-475.563$ 12466.46 0.2959 $-0.5827$ $-1.0$ 0.5558 119.406  108.1741 40
s5028 $-2157.179$ 599.608 $-396.271$ 10352.38 1.4017 $-0.9193$ $-1.0$ 2.9016 138.874   33.0037 28
s5030 $-2155.199$ 570.378 $-299.172$ 10324.29 1.6366 $-1.0334$ $-1.0$ 3.2711 131.218   30.0248 30
s5032 $-2172.510$ 583.512 $-348.992$ 10472.92 1.4649 $-0.9688$ $-1.0$ 2.9140 141.589   43.5871 32
$K\approx 250$ s5034 $-2179.610$ 584.069 $-331.053$ 10501.79 1.2593 $-0.9977$ $-1.0$ 2.5371 141.094   60.3202 34
s5036 $-2195.402$ 599.442 $-365.556$ 10609.25 0.9683 $-0.9629$ $-1.0$ 2.0020 145.638   80.1762 36
s5038 $-2205.358$ 602.093 $-375.998$ 10690.19 0.7284 $-0.9493$ $-1.0$ 1.5517 145.124   97.4925 38
s5040 $-2214.249$ 600.714 $-404.383$ 10802.96 0.5278 $-0.9019$ $-1.0$ 1.1569 146.358  112.2079 40
The density dependence of symmetry energy can be expanded around the saturation density $\rho_{0}$ as follows
$E_{\mathrm{sys}}(\rho)\approx{E_{\mathrm{sys}}(\rho_{0})}+\frac{L}{3}\left(\frac{\rho-\rho_{0}}{\rho_{0}}\right)+\frac{K_{\mathrm{sym}}}{18}\left(\frac{\rho-\rho_{0}}{\rho_{0}}\right)^{2}+\dots,$ (4)
where $L$ and $K_{\mathrm{sym}}$ are the slope and curvature of symmetry energy at nuclear saturation density $\rho_{0}$, respectively. These two quantities are defined as
$L$ $=$ $3\rho_{0}\frac{\partial{E_{\mathrm{sys}}(\rho)}}{{{\partial\rho}}}\Large{|}_{\rho=\rho_{0}},$ (5)
$K_{\mathrm{sym}}$ $=$ $9\rho_{0}^{2}\frac{\partial^{2}E_{\mathrm{sys}}(\rho)}{{\partial\rho^{2}}}\Large{|}_{\rho=\rho_{0}}.$ (6)
Also, the isoscalar incompressibility $K$ of symmetric nuclear matter is recalled as
$K=9\rho_{0}^{2}\frac{\partial^{2}E_{0}(\rho)}{{\partial\rho^{2}}}\Large{|}_{\rho=\rho_{0}}.$ (7)
Here, the quantity $E_{0}(\rho)$ is the energy per particle of symmetric nuclear matter Ding et al. (2023). The nuclear incompressibility can be extracted from the measurements of the isoscalar giant monopole resonance (ISGMR) in medium-heavy nuclei Colò et al. (2004); Cao et al. (2012). In addition, it is also mentioned that the incompressibility of symmetric nuclear matter can be deduced from the $\alpha$-decay properties Seif (2006). In Table 1, parameters of the Skyrme forces used in this work are shown. The corresponding saturation properties in infinite nuclear matter derived from the Skyrme forces, such as isoscalar incompressibility $K$, slope parameter $L$, and symmetry energy $E_{\mathrm{sys}}$, are also shown explicitly. Both the slope parameter $L$ and symmetry energy $E_{\mathrm{sys}}$ at saturation density cover a wide range.
III results and discussion
Linear correlation between the charge radii differences in the mirror partner nuclei 32Ar-32Si, 36Ca-36S, 38Ca-38Ar, and 54Ni-54Fe and the slope parameter of symmetry energy has been sequentially employed to constrain the equation of state of isospin asymmetric nuclear matter Brown et al. (2020); Pineda et al. (2021); König et al. (2023).
Table 2: $R_{\mathrm{ch}}$ and $\Delta{R_{\mathrm{ch}}}$ database for the $A=32$, $36$, $38$ and $54$ mirror-pair nuclei. The parentheses on the values of charge radii and the difference of charge radii are shown with systematic uncertainties Angeli and Marinova (2013); Miller et al. (2019); Pineda et al. (2021); Brown et al. (2020); König et al. (2023).
    $A$     $R_{\mathrm{ch}}$ (fm)     $\Delta{R_{\mathrm{ch}}}$ (fm)
32 Ar 3.3468(62)
Si 3.153(12) 0.194(14)
36 Ca 3.4484(27)
S 3.2982(12) 0.150(4)
38 Ca 3.4652(17)
Ar 3.4022(15) 0.063(3)
54 Ni 3.7370(30)
Fe 3.6880(17) 0.049(4)
In Table 2, the experimental data of $R_{\mathrm{ch}}$ and $\Delta{R_{\mathrm{ch}}}$ for $A=32$, $36$, $38$ and $54$ mirror-pair nuclei are given. As mentioned above, nuclear charge radius is influenced by the pairing correlations Reinhard and Nazarewicz (2017); An et al. (2020, 2022b). Thus the influence on the correlations between the charge radii difference in mirror partner nuclei and the slope parameter $L$ should be systematically investigated.
Figure 1: (Color online)Charge radii difference $\Delta{R_{\mathrm{ch}}}$ of mirror partner nuclei 54Ni-54Fe as a function of slope parameter $L$ is depicted under various pairing forces, e.g., no pairing (a), volume type $\eta=0.0$ (b), mixed type $\eta=0.5$ (c) and surface type $\eta=1.0$ (d) pairing forces. The Pearson coefficients $R$ are given in the parenthesis. The database takes from Refs. Angeli and Marinova (2013); Li et al. (2021); Pineda et al. (2021) and the symmetric error bands are labeled by light blue.
In Fig. 1, plots for charge radii difference $\Delta{R_{\mathrm{ch}}}$ in 54Ni-54Fe against symmetry energy slope $L$ are depicted by various pairing forces. For the convenience of discussion, the Pearson coefficient $R$, which is used to measure the degree of correlation between two variables, is employed in our study. Without considering the pairing correlations, the Pearson coefficient $R$ between $\Delta{R_{\mathrm{ch}}}$ of mirror partner nuclei 54Ni-54Fe and $L$ is given to be $0.993$. The same scenario can also be found in mixed-type pairing forces. For the volume-type pairing force, the Pearson coefficient $R$ is almost equivalent to no pairing and $\eta=0.0$ cases. As demonstrated in Ref. Huang et al. (2023), the linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$ is decreased by the pairing effects. From Fig. 1, one can find that the surface-type pairing force actually decrease the linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$ in mirror pair nuclei 54Ni-54Fe, in which the Pearson coefficient $R$ falls into $0.882$. Actually, more particles are scattered into the higher single levels under the surface pairing force. This leads to the rather poor correlation between the difference of charge radii of mirror partner nuclei and the slope parameter $L$.
Figure 2: (Color online)Charge radii differences $\Delta{R_{\mathrm{ch}}}$ of mirror partner nuclei 32Ar-32Si, 36Ca-36S, and 38Ca-38Ar as a function of slope parameter $L$ are depicted under no pairing and mixed-type $\eta=0.5$ pairing forces. The Pearson coefficients $R$ are given in the parenthesis. The database takes from Refs. Angeli and Marinova (2013); Li et al. (2021); Brown et al. (2020); König et al. (2023) and the symmetric error bands are labeled by light blue.
For further focusing on the linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$, charge radii differences $\Delta{R_{\mathrm{ch}}}$ of mirror partner nuclei 32Ar-32Si, 36Ca-36S, and 38Ca-38Ar against slope parameter $L$ are shown in Fig. 2. In this figure, clear linear correlations can be observed for these mirror-pair nuclei 32Ar-32Si, 36Ca-36S, and 38Ca-38Ar. For the volume-type pairing force, the Pearson coefficients are almost no change in the mirror partner nuclei 32Ar-32Si, 36Ca-36S and 38Ca-38Ar, namely the corresponding $R$ values are 0.926, 0.881 and 0.836, respectively. With considering the mixed-type pairing force, the Pearson coefficients $R$ are slightly improved with respect to the no pairing case. Actually, the Pearson coefficients $R$ are rather poor if the surface pairing force is taken into account. Here the corresponding figures are not shown, but the Pearson coefficients $R$ with considering surface-type pairing force are $0.887$, $0.791$ and $0.743$ for the corresponding mirror partner nuclei 32Ar-32Si, 36Ca-36S and 38Ca-38Ar, respectively. This means that pairing effect can decrease the linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$ if the pairing interactions are only tackled by the surface-type pairing force.
Figure 3: (Color online) Plots for the values of slope parameter $L$ extracted in this work (54Ni-54Fe, 38Ca-38Ar, 36Ca-36S, and 32Ar-32Si) and those from the existing literatures. The partially enumerated values of $L$, such as Wang $et~{}al$Wang and Li (2013), Brown $et~{}al$Brown et al. (2020), Pineda $et~{}al$Pineda et al. (2021), An $et~{}al$An et al. (2023c), Bano $et~{}al$Bano et al. (2023), König $et~{}al$König et al. (2023), pygmy dipole resonance (PDR) Carbone et al. (2010), electric dipole polarizability $\alpha_{\mathrm{D}}$ Roca-Maza et al. (2015), GW170817 Raithel and Özel (2019), pion production ($\pi^{-}/\pi^{+}$iu et al. (2023), and the charge exchange spin-dipole (SD) excitations Cheng et al. (2023), are also given. The color band represents the extracted slope parameter $L=42.57$-$50.64$ MeV in this literature.
As mentioned in Refs. Wang and Li (2013); Brown (2017); Brown et al. (2020); Pineda et al. (2021); König et al. (2023), charge radii differences in mirror partner nuclei provide an alternative approach to pin down the isovector components in the equation of state of asymmetric nuclear matter. To facilitate the relatively clear linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$, the mixed-type pairing force is employed in the following discussion. In Fig. 3, the deduced slope parameters $L$ from the differences of charge radii of mirror-pair nuclei 54Ni-54Fe, 38Ca-38Ar, 36Ca-36S, and 32Ar-32Si are drawn by shaping the covered range in Fig. 2. Combing the experimental data, the extracted slope parameter falls into the range of $L=42.57$-$50.64$ MeV. From this figure, one can find that the deduced value of $L$ covers the uncertainty range deriving from the mirror charge radii differences Wang and Li (2013); Brown et al. (2020); Pineda et al. (2021); An et al. (2023c); Bano et al. (2023); König et al. (2023). Here it is mentioned that the upper limit range of $L$ coming from 32Ar-32Si is 96.65 MeV. This result is overestimated in comparison with the Ref. König et al. (2023) where the upper range is confined to be $L\leq 60$ MeV. This may be due to the fact that shape deformation is ignored in our study. Besides, the pairing correlations and the limited Skyrme parameter sets used here may also have an influence on evaluating the interval range of slope parameter $L$.
Also, the value of symmetry energy slope $L$ extracted from pygmy dipole resonance (PDR) locates at the internal range of $L=64.8\pm 15.7$ Carbone et al. (2010). The covered range is very narrow with respect to other cases of this figure. Furthermore, the extracted $L$ in this work can cover the uncertainty ranges deriving from the nuclear electric dipole polarizability $\alpha_{\mathrm{D}}$ Roca-Maza et al. (2015), dense astrophysical event GW170817 Raithel and Özel (2019), the pion production ratios in heavy-ion collisions Liu et al. (2023) and the charge exchange spin-dipole (SD) excitations Cheng et al. (2023). Besides, this deduced result can mostly cover the range of $L$ shown in Ref. Zhang and Chen (2014), in which the slope parameter captures the range of $L=47.3\pm 7.8$. In a word, the extracted range of slope parameter $L=42.57$-$50.64$ MeV means that the rather soft equation of state of isospin asymmetric nuclear matter can be obtained from the implication of charge radii difference in mirror-pair nuclei.
Figure 4: (Color online) Charge radii differences $\Delta{R_{\mathrm{ch}}}$ of mirror partner nuclei 36Ca-36S and 38Ca-38Ar are plotted as a function of slope parameter $L$ classified by various isoscalar incompressibility coefficients $K$. The corresponding experimental data are taken from Ref. Brown et al. (2020).
As shown in Fig. 2, the Pearson coefficients become lower for the mirror-pair nuclei 36Ca-36S and 38Ca-38Ar. This suggests that the profoundly theoretical uncertainties are encountered in the calibrating protocol. In order to clarify this phenomenon, the charge radii differences $\Delta{R_{\mathrm{ch}}}$ of mirror partner nuclei 36Ca-36S and 38Ca-38Ar against the slope parameter $L$ classified by various isoscalar incompressibility coefficients $K$ are shown in Fig. 4. From both panels of Fig. 4, one can find that the isoscalar incompressibility has an influence on the charge radii difference of mirror partner nuclei. Even from $K\approx 230$ to $K\approx 250$ MeV, the deviated range for $L$ is about $38.0$ MeV. Here, as shown in Fig. 3, we mention that the influence of the isoscalar incompressibility coefficient cannot be taken into account in evaluating the slope parameter $L$. Significantly the isoscalar incompressibility may be indispensable in calibrating protocol. As is well known, the accurate description of the density-dependent behavior of symmetry energy plays an important role in recognizing various physical mechanisms through nuclear models, such as the production of superheavy elements Veselsky et al. (2016), and the cluster radioactivity Xu et al. (2014); Yang et al. (2023b), etc. Especially, recent study suggests that the effective proton-neutron chemical potential difference of neutron-rich nuclei is strongly sensitive to the symmetry energy Qiu et al. (2023). Therefore, more aspects should be taken into account in constraining the slope parameter $L$.
IV summary and outlook
In this work, the correlation between the charge radii differences ($\Delta{R_{\mathrm{ch}}}$) in mirror-pair nuclei 32Ar-32Si, 36Ca-36S, 38Ca-38Ar, and 54Ni-54Fe and slope parameter $L$ at saturation density is investigated by the spherical Skyrme energy density functionals. The calculated results suggest that the surface pairing force can further decrease the linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$. The deduced slope parameter locates the interval range of $L=42.57$-$50.64$ MeV under the mixed pairing interaction, namely the rather soft equation of state. This is in accord with most existing literatures. Meanwhile, the influence of isoscalar nuclear matter property on $\Delta{R_{\mathrm{ch}}}$ is inevitable in constraining the nuclear matter EoS. This may suggest that most parametrization forces are ruled out in the calibrated procedure. As shown in Fig. 4, the correlation between charge radii difference of mirror nuclei and the slope parameter of symmetry energy is influenced by nuclear incompressibility. This seems to suggest that the Skyrme parameter sets adjusted to different slope parameters under specific incompressibility coefficient can give highly linear correlation between the charge radii difference of mirror partner nuclei and the slope parameter $L$. Meanwhile, this may provide an alternative approach to evaluate the possible influential quantities in determining the EoS of asymmetric nuclear matter. Shown in Ref. Huang et al. (2023), the calculated results obtained by density energy functionals consisting of Skyrme and covariant models suggest that more mirror-pair nuclei are ruled out in constraining the slope parameter $L$, but the superior candidates of mirror-pair nuclei 44Cr-44Ca and 46Fe-46Ca may be used to ascertain the limit range of $L$. Thus more discussion should be performed in the proceeding work.
Recently studies show that the charge-changing cross section measurements can be used to determine the EoS of nuclear matter with mirror nuclei Xu et al. (2022b). This method can also be employed to investigate the charge radii of exotic nuclei Wang et al. (2023) and the isospin interactions Zhao et al. (2023). This provides an accessable approach to ascertain the nuclear symmetry energy. From the aspect of the charge radii difference in mirror-pair nuclei, the underlying mechanism should be taken into account in determining the nuclear symmetry energy, such as the local relations among the adjacent neighbouring nuclei Sun et al. (2014); Bao et al. (2016, 2020); Ma et al. (2021), the $\alpha$-cluster formation Qian et al. (2014); Ni et al. (2013), etc. Thus more high-precision charge radii data are urgently required in experimental and theoretical researches.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China under Grants No. 12275025, No. 11975096 and the Fundamental Research Funds for the Central Universities (2020NTST06).